OpenStudy (anonymous):

find the surface area of a rectangular prism with the length 18 inches, width 6 inches, and height 4 inches. A. 256in^2 B 204 in^2 C. 507 in^2 D. 408 in^2

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I got 408

OpenStudy (firejay5):

If you are having trouble with finding the surface area or volume google this stuff

OpenStudy (anonymous):

@katlin95 use the formula @jim_thompson5910 gave you before in your other question.

jimthompson5910 (jim_thompson5910):

that was for volume, use this one for surface area SA = 2*(LW + WH + LH)

OpenStudy (anonymous):

my computer is being slow, so its hard for me to go back and fourth

OpenStudy (anonymous):

I got C

jimthompson5910 (jim_thompson5910):

use SA = 2*(LW + WH + LH) where L = 18 W = 6 H = 4

OpenStudy (firejay5):

It's 408

OpenStudy (anonymous):

lol u shoulda told me I was right then lol

OpenStudy (firejay5):

If you google this information it wouldn't be hard

jimthompson5910 (jim_thompson5910):

sry i didn't see the 408 you posted katlin95, but you're right